OpenStudy (anonymous):

I really need help with systems of equations! 3x+4y=8 -2x+5y=3 (I need to find x and y)

hero (hero):

First add both equations together. Let me know what you get afterwards.

OpenStudy (anonymous):

what do you mean add them?

OpenStudy (anonymous):

oh I got it! ok one sec

OpenStudy (anonymous):

-5x+9y=11?

hero (hero):

You added 3 and -2 and got 5?

hero (hero):

Doesn't 3 + (-2) = 1

OpenStudy (anonymous):

yeah yeah, sorry. so 1x+9y=11?

hero (hero):

Yes or simply x + 9y = 11

hero (hero):

Now, with this form, what we can do next is isolate x by subtracting 9y from both sides to get x = 11 - 9y

OpenStudy (anonymous):

ok :)

hero (hero):

Now we can just simply substitute the expression 11 - 9y for x in to one of the original equations.

hero (hero):

3(11 - 9y) + 4y = 8

hero (hero):

At this point, you can solve for y

OpenStudy (anonymous):

y=29/27?

hero (hero):

Can you show your work please?

OpenStudy (anonymous):

actually I did it with a calculator haha sorry, I didnt want to seem dumb. I´ll try though

hero (hero):

You got y = 29/27 from a calculator? Are you sure?

OpenStudy (anonymous):

yeah, an online calculator

OpenStudy (anonymous):

anyways Ill try to do it step by step

hero (hero):

Either that calculator isn't working properly or you input the equation incorrectly.

OpenStudy (anonymous):

3(11-9y)=8-4y

OpenStudy (anonymous):

(11-9y)=4/3y?

hero (hero):

Not exactly. Why would you divide both sides by 3 instead of distributing the 3?

hero (hero):

Do you remember distributive property? a( b + c) = ab + ac 3(11 - 9y) = ?

OpenStudy (anonymous):

3(11-9y)=4?

OpenStudy (anonymous):

thank you for your patience, I am the worst at math I know

OpenStudy (anonymous):

y=25/23 ? @Hero

hero (hero):

y = 25/23 but I'm more interested in the steps you took to get that.

OpenStudy (anonymous):

I thought you left, thats why I just wanted to quickly check with you if it was the right answer before you left

OpenStudy (anonymous):

so 3(11-9y)=4?

hero (hero):

It is not possible to remove the y from the 4 without justification. I don't understand what you believe you're trying to accomplish here. What happened to distributing the 3?

OpenStudy (anonymous):

I dont understand either haha. ok so if we have 3(11-9y)+4=8, what would be the next step?

hero (hero):

The original equation was 3x + 4y = 8 correct?

OpenStudy (anonymous):

yes.

hero (hero):

And we already solved for x and got 11 - 9y correct?

OpenStudy (anonymous):

yes.

hero (hero):

So if we remove x from 3x + 4y = 8 and replace it with 11 - 9y, we get 3(11 - 9y) + 4y = 8

hero (hero):

The next step is to distribute the 3

OpenStudy (anonymous):

thats the part where i am stuck, how do I distribute it?

hero (hero):

3(11 - 9y) = 33 - 27y

OpenStudy (anonymous):

ok, so 3(11-9y)=6y?

OpenStudy (anonymous):

@hero?